# Exact WKB in all sectors I: Potentials with degenerate saddles

### v1

@@ 2.2 Weber-type approach
So far, in the literature, the Weber-type EWKB is a well-discussed with a focus on the region below the barrier top Sueishi:2021xti ; Kamata:2021jrs ; Kamata:2023opn ; Kamata:2024tyb . In these cases, the information about the WKB cycles arises from the minima of the potential: The perturbative part is given by a residue integration around the saddle point. The non-perturbative part, on the other hand, is more intricate; but it is possible to obtain a generic expression. It can be shown that it has a generic structure which can be separated into two parts: One originates from the locally harmonic saddle point and carries the local information. The other one arises from the connection of successive saddle points and can be identified as a non-local expression.
@@
In this subsection, we discuss the generalization of the Weber-type EWKB analysis of Sueishi:2021xti to all types of locally harmonic saddles, i.e. both minimum and maximum on the real line, of a given 1D potential. Our main purpose is obtaining the generic expressions for perturbative and non-perturbative WKB actions around both the bottom of wells and the top of barriers. Previously, this was done for non-perturbative cycles connecting two degenerate minima in Sueishi:2021xti , where the non-perturbative action was expressed as [^5] The notation will be clarified later in this section.
$e^{\frac{2\pi i}{g}a^{\mathrm{D}}(\tilde{u},g)}=\mathcal{K}\,\frac{2\pi}{\left(\Gamma\left(\frac{1}{2}+\mathcal{F}(\tilde{u},g)\right)\right)^{2}}\,g^{-2\mathcal{F}(\tilde{u},g)}.$ (27)
While generalizing this discussion to all types of locally harmonic saddles, we also handle the Weber-type EWKB analysis in a slightly different way. Specifically, we keep the normalization of the formal WKB solutions at the turning points rather than setting it to infinity as in Sueishi:2021xti . We observe that our setup simplifies the analysis of the non-perturbative cycles and differentiates the local and non-local parts in a more explicit manner.
@@
In addition to this general construction, we also provide generic expressions of the $\mathcal{K}$ factor in (27) under general assumptions about the turning points that the Weber-type WKB cycles connect. Note that for the saddles at the bottom of the wells, this generic expression was first obtained by Zinn-Justin and Jentschura in relation with the multi-instanton (both bions and bounces) configurations Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw . In this section, we recover them by the Weber-type EWKB and generalize them to all saddles of a given potential.
@@
As the name suggests, the Weber-type EWKB is originated from the Weber equation. For our purposes, it represents the local behaviour around saddle points and its EWKB analysis is linked to the EWKB analysis of the locally harmonic potentials via local-global relations. The non-local quantities, on the other hand, will arise from the connection of successive saddle points, which is related to change of the normalization of the formal solution from EWKB point of view.
@@
    $\left(-g^{2}\frac{{\mathrm{d}}^{2}}{{\mathrm{d}}y^{2}}+Q(y,\tilde{u})\right)\Psi_{\mathrm{L}}(y))=0,$ (28)
@@
Higher order terms can be found recursively if needed. Note that for a given potential $V(x)$, the WKB expansions of $\tilde{S}$ in the Weber type and $\tilde{s}$ in the Airy type are equal to each other as $\tilde{S}$ can be obtained by expanding $\tilde{s}$ for around $u=u_{0}$.
@@
Using this series expansion, (39) can be solved for $y_{m}$ order by order in $g$ which becomes very useful in practical calculations.
@@
Finally, following the arguments in Kawai1 , we can show that the Stokes discontinuities of local and global solutions are the same. Therefore, it is enough to obtain the monodromy matrices in the local coordinates and the same matrices applies in the global coordinates provided by appropriate coordinate transformations. This equivalence is the source of the “universal” behavior of the Stokes discontinuities for generic potentials in global coordinates as noticed before in Kawai1 ; Bucciotti:2023trp . As we will see later, the Stokes discontinuities determine the local structure of non-perturbative cycles which also contains a non-local part. Despite this additional part, we observe that non-perturbative cycles also admit a “universal” structure.
@@
Let us first focus on the local part and perform a EWKB analysis on the Weber equation. In the following, we briefly review the construction of the Stokes geometry for the Weber equation with generic $\omega$. As it was discussed in detail in Sueishi:2021xti ; Kamata:2021jrs ; Kamata:2023opn ; Kamata:2024tyb for $\omega^{2}=1$, we mainly emphasize the generalization to generic $\omega$.
@@
where the integral is $\Psi_{\mathrm{L}}^{-}$ in (46) for $y=|y|\,e^{i\theta_{y}}$ and the discontinuity of $\Psi^{+}$ is written as
@@
Other Stokes lines, $\arg(\omega y^{2})=\pm\pi$ and $\arg(\omega y^{2})=2\pi$, lie on the directions $\theta_{y}\pm\frac{\pi}{2}$ and $\theta_{y}+\pi$, respectively. The discontinuities on these directions are found as
@@
Note that the prefactors on RHS of (52)-(55) have the same structure up to some signs and a phase which is governed by $\arg(\omega y^{2})$. Setting $\theta=\arg(\omega y^{2})$, we can write the discontinuities in a compact form as
$\Delta\Psi^{+}_{\mathrm{L},\theta}=e^{-\frac{i\theta\tilde{u}}{\omega}}\,\Phi^{+}_{\omega}(\tilde{u},g)\,\Psi^{-}_{\mathrm{L},\theta}\,,\qquad\qquad\Delta\Psi^{-}_{\mathrm{L},\theta}=e^{\frac{i\theta\tilde{u}}{\omega}}\,\Phi^{-}_{\omega}(\tilde{u},g)\,\Psi^{+}_{\mathrm{L},\theta},$ (56)
@@
(a) Diagrams when $\arg\omega=0$ and $\arg\omega=\pi$
(b) Diagrams when $\arg\omega=\pm\frac{\pi}{2}$
Figure 6: Stokes Diagrams of the Weber equation for $\omega^{2}=\pm 1$
@@
$\theta_{y}$ = 0 $\Delta\Psi^{+}_{\mathrm{L}}=\Phi_{\omega=1}^{+}\,\Psi^{-}_{\mathrm{L}}$ $\theta_{y}=0$ $\Delta\Psi^{-}_{\mathrm{L}}=e^{-i\pi\tilde{u}}\,\Phi^{-}_{\omega=-1}\,\Psi^{+}_{\mathrm{L}}$
$\theta_{y}$ = $\frac{\pi}{2}$ $\Delta\Psi^{-}_{\mathrm{L}}=e^{i\pi\tilde{u}}\;\Phi_{\omega=1}^{-}\,\Psi^{+}_{\mathrm{L}}$ $\theta_{y}=\frac{\pi}{2}$ $\Delta\Psi^{+}_{\mathrm{L}}=e^{2\pi i\tilde{u}}\,\Phi^{+}_{\omega=-1}\,\Psi^{-}_{\mathrm{L}}$
$\theta_{y}$ = $\pi$ $\Delta\Psi^{+}_{\mathrm{L}}=e^{-2i\pi\tilde{u}}\;\Phi_{\omega=1}^{+}\,\Psi^{-}_{\mathrm{L}}$ $\theta_{y}=-\pi$ $\Delta\Psi^{-}_{\mathrm{L}}=e^{i\pi\tilde{u}}\,\Phi F^{-}_{\omega=-1}\,\Psi^{+}_{\mathrm{L}}$
@@
In addition to the new results in Table 2.2, in Table 2.2 we present the results for $\arg\omega=0$ which recover the discontinuities in Sueishi:2021xti ; Bucciotti:2023trp . For $\arg\omega=\pi$, on the other hand, the discontinuities slightly differ from Sueishi:2021xti , where the associated Stokes diagram was considered without the change in $\arg\omega$ being taken into account. These differences, however, are merely related to our convention on Stokes diagrams and it doesn’t affect the final result when Weber-type EWKB is discussed in global coordinates.
@@
Now, we can construct the contribution of WKB cycles in the Weber-type approach for possible Stokes diagrams of generic locally harmonic oscillators. In Fig. 7, we present all possible cases of the Weber-type Stokes diagrams around the bottom of wells and the top of barriers. We again emphasize that the transition to the Weber-type Stokes diagrams is only related to the scaling of the spectral parameter, i.e. $u\rightarrow g\tilde{u}$ and rescaling of the potential, i.e. $V(x)\rightarrow V(x)-u_{0}$. As a result of this, the Airy-type Stokes diagrams for $\theta_{u}=0$ and $\theta_{u}=\pm\pi$ in Fig. 5 turn into the Weber-type Stokes diagrams around minima and around maxima respectively as depicted in Figures 6(a) and 6(b).
(c) 2 saddle points
(d) 1 saddle point and 1 simple turning point
@@
is the matrix for changing the normalization point and $M^{+}_{\mathrm{A}}$ is the monodromy matrix for Airy type diagrams originated from the simple turning point at $x=x_{0}$ in Fig. 6(d) . Then, we get the associated non-perturbative actions as
@@
More specifically, if we focus on the bottom of wells, in the case of Fig. 6(c), ${\mathcal{B}}$ contributes to $\Pi_{\mathrm{NP}}^{\mathrm{i}}$ as the (neutral) bion action which is a configuration of an instanton and an anti-instanton; while, in the case of Fig. 6(d), ${\mathcal{B}}$ contributes to $\Pi_{\mathrm{NP}}^{\mathrm{ii}}$ as a bounce moving on a closed path between a saddle point, i.e. $x_{1}$ in this case, and a simple turning point, i.e. $x_{0}$.
@@
The computations of higher order corrections to $W_{0,1}$, on the other hand, are more involved. For example, at the next order, we encounter the following integral, when $x_{0}$ and $x_{1}$ are double saddle points as in Fig. 6(c):
@@
$\simeq\frac{\tilde{u}}{\omega_{0}}\Big{[}\log\tilde{\mathcal{C}}_{\mathrm{i}}-\log\left(x_{1}-x\right)\Big{]}_{x=x_{1}}-\frac{\tilde{u}}{\omega_{0}}\Big{[}-\log\tilde{\mathcal{C}}_{\mathrm{i}}+\log\left(x-x_{0}\right)\Big{]}_{x=x_{0}}\,.$ (73)
When only $x_{1}$ is the double turning point as in Fig. 6(d), we have
$\mathcal{K}_{\mathrm{ii}}$ $=\int_{x_{0}}^{x_{1}}{\mathrm{d}}x\,\frac{\tilde{u}}{\omega_{0}}\frac{\omega_{0}}{\sqrt{2\tilde{V}(x)}}\simeq\frac{\tilde{u}}{\omega_{0}}\Big{[}\log\mathcal{C}_{\mathrm{ii}}-\log\left(x_{1}-x\right)\Big{]}_{x=x_{1}}\,.$ (74)
In both cases, $\tilde{\mathcal{C}}_{\mathrm{i}}$ and $\tilde{\mathcal{C}}_{\mathrm{ii}}$ are constants specified by the potential $\tilde{V}(x)$.
Note that in (73) and (74), we inserted $\omega_{0}$ on purpose since $\frac{\tilde{u}}{\omega_{0}}$ is the term appearing in $\mathcal{F}_{\omega_{0}}$. This property can be deduced from the residue computation in the local coordinates in (58) which turns into the leading order term of $\mathcal{F}_{\omega_{0}}(\tilde{u},g)$ at $g=0$. In fact, at any order contributing to (71), we encounter such integrals which generate the corrections to $\log\tilde{\mathcal{C}}_{\mathrm{i}}$ and $\log\tilde{\mathcal{C}}_{\mathrm{ii}}$ terms in (73) and (74) respectively. Eventually, their total contribution to $e^{-W_{0,1}}$ becomes
$e^{-W_{0,1}}=e^{-\mathcal{G}(\tilde{u},g)}e^{-2\mathcal{F}(\tilde{u},g)\log\tilde{\mathcal{C}}_{\mathrm{i}}}\,,$ (75)
@@
$e^{-W_{0,1}}=e^{-\mathcal{G}(\tilde{u},g)}e^{-\mathcal{F}(\tilde{u},g)\log\tilde{\mathcal{C}}_{\mathrm{ii}}}\,,$ (76)
@@
In this form, the computations of $W_{0,1}$ can be carried out by using Mellin transformation at each order as described in Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw ; Kamata:2021jrs . Then, following Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw , we introduce normalization at the double turning points and dropping $\omega_{0}$ subscripts. We write the final generic expressions for $\Pi_{\mathrm{NP}}^{\mathrm{i}}$ and $\Pi_{\mathrm{NP}}^{\mathrm{ii}}$ as
$\tilde{\Pi}^{\mathrm{i}}_{\mathrm{NP}}$ $=e^{-\mathcal{G}(\tilde{u},g)}\,\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{i}}}\right)^{-2\mathcal{F}(\tilde{u},g)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}(\tilde{u},g)\right)\right]^{2}}\,,\qquad\tilde{\Pi}^{\mathrm{ii}}_{\mathrm{NP}}=e^{-\mathcal{G}(\tilde{u},g)}\,\frac{\sqrt{2\pi}\,\left(\frac{g}{\mathcal{C}_{\mathrm{ii}}}\right)^{-\mathcal{F}(\tilde{u},g)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}(\tilde{u},g)\right)}\,,$ (79)
@@
$\mathcal{C}_{\mathrm{i}}$ $=2\,\omega_{0}\left(x_{1}-x_{0}\right)^{2}\,\mathrm{exp}\left\{\int_{x_{0}}^{x_{1}}{\mathrm{d}}x\,\left[\frac{\omega_{0}}{\sqrt{2V(x)}}-\frac{1}{x-x_{0}}-\frac{1}{x_{1}-x}\right]\right\}\,,$ (80)
$\mathcal{C}_{\mathrm{ii}}$ $=2\,\omega_{0}\,\left(x_{1}-x_{0}\right)^{2}\mathrm{exp}\left\{2\int_{x_{0}}^{x_{1}}{\mathrm{d}}x\,\left[\frac{\omega_{0}}{\sqrt{2V(x)}}-\frac{1}{x_{1}-x}\right]\right\}\,.$ (81)
Note that we inserted the prefactor $\omega_{0}$, which originally appears in $\frac{g}{\omega_{0}}$ terms in (70) in the definitions of $\mathcal{C}_{\mathrm{i}}$ and $\mathcal{C}_{\mathrm{ii}}$ for convenience.
These expressions recover Zinn-Justin’s results in Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw for the bottom of wells and generalizes them for all types of locally harmonic saddles. At the barrier tops, $\omega_{0}$ is an imaginary constant and another imaginary contribution arises from $\sqrt{2V(x)}$. As a result, the term $\frac{\omega_{0}}{\sqrt{V(x)}}$ is real and its log divergence(s) at double turning point(s) is(are) cancelled by the normalization terms in (80) and (81).

### v2

@@ 2.2 Weber-type approach
So far, in the literature, the analysis of one dimensional quantum mechanics in view of the Weber-type EWKB is well-discussed with a focus on the region below the barrier top Sueishi:2021xti ; Kamata:2021jrs ; Kamata:2023opn ; Kamata:2024tyb . In these cases, the information about the WKB cycles arises from the minima of the potential: The perturbative part is given by a residue integration around the saddle point. The non-perturbative part, on the other hand, is more intricate; but it is possible to obtain a generic expression, which can be separated into two parts: One originates from the locally harmonic saddle point and carries the local information. The other one arises from the connection of successive turning points and can be identified as a non-local expression.
The analysis of barrier top region with the Weber-type EWKB, on the other hand, is much less discussed in the literature. In highly sophisticated mathematical papers DDP2 ; DP1 , Delabaere, Dillinger and Pham investigated the Weber-type EWKB around all type of locally harmonic saddles including the barrier tops. In this subsection, we approach the problem from a different point of view [^5] This difference was also emphasized in DP1 where the authors call their method uses the transformations of resurgent functions while the constructions in AKT1 ; AKT4 are based on micro-differential analysis which we utilize in this paper. based on AKT1 ; AKT4 (see also Kawai1 ), which was applied to the quantum mechanical problems in Sueishi:2021xti ; Kamata:2021jrs . We also provide a construction from a physicist point of view with a connection to the relevant physics literature.
@@
In the following, we follow the construction in Sueishi:2021xti and generalize it to all types of locally harmonic saddles, i.e. both minimum and maximum on the real line, of a given 1D potential. Our main purpose is obtaining the generic expressions for perturbative and non-perturbative WKB actions around both the bottom of wells and the top of barriers. Previously, this was done for non-perturbative cycles connecting two degenerate minima in Sueishi:2021xti , where the non-perturbative action was expressed as [^6] The notation will be clarified later in this section.
$e^{\frac{2\pi i}{g}a^{\mathrm{D}}(\tilde{u},g)}=\mathcal{K}(\tilde{u},g)\,\frac{2\pi}{\left(\Gamma\left(\frac{1}{2}+\mathcal{F}(\tilde{u},g)\right)\right)^{2}}\,g^{-2\mathcal{F}(\tilde{u},g)}.$ (27)
Note that in DP1 , the application of Weber-type EWKB to the maxima of double-well potential was discussed where the authors computed $\mathcal{F}$ and $\mathcal{K}$ order by order in $g$ for the double-well potential and inferred the prefactor $\mathcal{K}$ from these expansions.
@@
In our discussion, on the other hand, we provide generic expressions of the $\mathcal{K}$ factor in (27) under general assumptions about the turning points that the Weber-type WKB cycles connect. Note that for the saddles at the bottom of the wells, it was first obtained by Zinn-Justin and Jentschura (ZJJ) in relation with the multi-instanton (both bions and bounces) configurations Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw ; Jentschura:2010zza . In this section, we recover them by the Weber-type EWKB and generalize them to all saddles of a given potential. With this generalization, we observe the exchange of the perturbative and non-perturbative cycles around the maxima of the potential in comparison to its minima. This manifestly points out an $S$-duality between minima and maxima of a classical potential from the EWKB point of view.
Finally, note that in the following discussion, we also handle the Weber-type EWKB analysis in a slightly different way than Sueishi:2021xti . Specifically, we keep the normalization of the formal WKB solutions at the turning points rather than setting it to infinity as in Sueishi:2021xti . We observe that our setup simplifies the analysis of the non-perturbative cycles and differentiates the local and non-local parts in a more explicit manner.
@@
As the name suggests, the Weber-type EWKB is originated from the Weber equation. For our purposes, it represents the local behaviour around saddle points and its EWKB analysis is linked to the EWKB analysis of the locally harmonic potentials via local-global relations.
@@
    $\left(-g^{2}\frac{{\mathrm{d}}^{2}}{{\mathrm{d}}y^{2}}+Q(y,\tilde{u})\right)\Psi_{\mathrm{L}}(y)=0,$ (28)
@@
Higher order terms can be found recursively if needed. Note that for a given potential $V(x)$, the expansions of $\tilde{S}$ in the Weber type EWKB and $\tilde{s}$ in the Airy type EWKB are equal to each other as $\tilde{S}$ can be obtained by expanding $\tilde{s}$ for around $u=u_{0}$.
@@
Finally, following the arguments in Kawai1 , we can show that the Stokes discontinuities of local and global solutions are the same. Therefore, it is enough to obtain the monodromy matrices in the local coordinates and the same matrices applies in the global coordinates provided by appropriate coordinate transformations. This equivalence is the source of the “universal” behavior of the Stokes discontinuities for generic potentials in global coordinates as noticed before in DDP2 ; DP1 ; Kawai1 ; Bucciotti:2023trp . Note that as we will see later, the Stokes discontinuities determine the local structure of non-perturbative cycles which also contains a non-local part. Despite this additional part, we observe that non-perturbative cycles also admit a “universal” structure.
@@
Let us first focus on the local part and perform a EWKB analysis on the Weber equation. In the following, we briefly review the construction of the Stokes geometry for the Weber equation with generic $\omega$. This was discussed in detail in Sueishi:2021xti ; Kamata:2021jrs ; Bucciotti:2023trp for $\omega^{2}=1$. Here, we mainly emphasize the generalization to generic $\omega$.
@@
where the integral is $\Psi_{\mathrm{L}}^{-}$ in (46) for $y=|y|\,e^{i\theta_{y}}$. Then, the discontinuity of $\Psi^{+}_{\mathrm{L},\theta_{y}}$ is written as
@@
Other Stokes lines, $\arg(\omega y^{2})=\pm\pi$ and $\arg(\omega y^{2})=2\pi$, lie on the directions $\theta_{y}\pm\frac{\pi}{2}$ and $\theta_{y}+\pi$, respectively. The discontinuities on these directions are written as
@@
Note that the prefactors on RHS of (52)-(55) have the same structure up to some signs and a phase which is governed by $\arg(\omega y^{2})$. Setting $\alpha=\arg(\omega y^{2})$, we can write the discontinuities in a compact form as
$\Delta\Psi^{+}_{\mathrm{L},\theta}=e^{-\frac{i\alpha\tilde{u}}{\omega}}\,\Phi^{+}_{\omega}(\tilde{u},g)\,\Psi^{-}_{\mathrm{L},\theta}\,,\qquad\qquad\Delta\Psi^{-}_{\mathrm{L},\theta}=e^{\frac{i\alpha\tilde{u}}{\omega}}\,\Phi^{-}_{\omega}(\tilde{u},g)\,\Psi^{+}_{\mathrm{L},\theta},$ (56)
@@
In addition to the new results in Table 2.2, we present the results for $\arg\omega=0$ in Table 2.2, which recover the discontinuities in Sueishi:2021xti ; Bucciotti:2023trp . For $\arg\omega=\pi$, on the other hand, the discontinuities slightly differ from Sueishi:2021xti , where the associated Stokes diagram was considered without the change in $\arg\omega$ being taken into account. These differences, however, are merely related to our convention on Stokes diagrams and it doesn’t affect the final result when Weber-type EWKB is discussed in global coordinates.
@@
$\theta_{y}$ = 0 $\Delta\Psi^{+}_{\mathrm{L}}=\Phi_{\omega=1}^{+}\,\Psi^{-}_{\mathrm{L}}$ $\theta_{y}=0$ $\Delta\Psi^{-}_{\mathrm{L}}=e^{i\pi\tilde{u}}\,\Phi^{-}_{\omega=-1}\,\Psi^{+}_{\mathrm{L}}$
$\theta_{y}$ = $\frac{\pi}{2}$ $\Delta\Psi^{-}_{\mathrm{L}}=e^{i\pi\tilde{u}}\;\Phi_{\omega=1}^{-}\,\Psi^{+}_{\mathrm{L}}$ $\theta_{y}=\frac{\pi}{2}$ $\Delta\Psi^{+}_{\mathrm{L}}=e^{-2\pi i\tilde{u}}\,\Phi^{+}_{\omega=-1}\,\Psi^{-}_{\mathrm{L}}$
$\theta_{y}$ = $\pi$ $\Delta\Psi^{+}_{\mathrm{L}}=e^{-2i\pi\tilde{u}}\;\Phi_{\omega=1}^{+}\,\Psi^{-}_{\mathrm{L}}$ $\theta_{y}=-\pi$ $\Delta\Psi^{-}_{\mathrm{L}}=e^{-i\pi\tilde{u}}\,\Phi^{-}_{\omega=-1}\,\Psi^{+}_{\mathrm{L}}$
@@
(a) Diagrams when $\arg\omega=0$ and $\arg\omega=\pi$
(b) Diagrams when $\arg\omega=\pm\frac{\pi}{2}$
Figure 6: Stokes Diagrams of the Weber equation for $\omega^{2}=\pm 1$
@@
Now, we can construct the contribution of WKB cycles in the Weber-type approach for possible Stokes diagrams of generic locally harmonic oscillators. In Fig. 7, we present all possible cases of the Weber-type Stokes diagrams around the bottom of wells and the top of barriers. We again emphasize that the transition to the Weber-type Stokes diagrams is only related to the scaling of the spectral parameter, i.e. $u\rightarrow g\tilde{u}$ and rescaling of the potential, i.e. $V(x)\rightarrow V(x)-u_{0}$. As a result of this, the Airy-type Stokes diagrams for $\theta_{u}=0$ and $\theta_{u}=\pm\pi$ in Fig. 5 turn into the Weber-type Stokes diagrams around minima and around maxima respectively as depicted in Figures 5(a) and 5(b).
(a) 2 saddle points
(b) 1 saddle point and 1 simple turning point
@@
is the matrix for changing the normalization point and $M^{+}_{\mathrm{A}}$ is the monodromy matrix for Airy type diagrams originated from the simple turning point at $x=x_{0}$ in Fig. 6(b) . Then, we get the associated non-perturbative actions as
@@
More specifically, if we focus on the bottom of wells, in the case of Fig. 6(a), ${\mathcal{B}}$ contributes to $\Pi_{\mathrm{NP}}^{\mathrm{i}}$ as the (neutral) bion action which is a configuration of an instanton and an anti-instanton; while, in the case of Fig. 6(b), ${\mathcal{B}}$ contributes to $\Pi_{\mathrm{NP}}^{\mathrm{ii}}$ as a bounce moving on a closed path between a saddle point, i.e. $x_{1}$ in this case, and a simple turning point, i.e. $x_{0}$.
@@
The computations of higher order corrections to $W_{0,1}$, on the other hand, are more involved. For example, at the next order, we encounter the following integral, when $x_{0}$ and $x_{1}$ are double saddle points as in Fig. 6(a):
@@
$\simeq\frac{\tilde{u}}{\omega_{0}}\Big{[}\log\mathcal{C}^{\mathrm{i}}_{0}-\log\left(x_{1}-x\right)\Big{]}_{x=x_{1}}-\frac{\tilde{u}}{\omega_{0}}\Big{[}-\log\mathcal{C}^{\mathrm{i}}_{0}+\log\left(x-x_{0}\right)\Big{]}_{x=x_{0}}\,.$ (73)
When only $x_{1}$ is the double turning point as in Fig. 6(b), we have
$\mathcal{K}_{\mathrm{ii}}$ $=\int_{x_{0}}^{x_{1}}{\mathrm{d}}x\,\frac{\tilde{u}}{\omega_{0}}\frac{\omega_{0}}{\sqrt{2\tilde{V}(x)}}\simeq\frac{\tilde{u}}{\omega_{0}}\Big{[}\log\mathcal{C}^{\mathrm{ii}}_{0}-\log\left(x_{1}-x\right)\Big{]}_{x=x_{1}}\,.$ (74)
In both cases, $\mathcal{C}^{\mathrm{i}}_{0}$ and $\mathcal{C}^{\mathrm{ii}}_{0}$ are constants specified by the potential $\tilde{V}(x)$.
Note that in (73) and (74), we inserted $\omega_{0}$ on purpose since $\frac{\tilde{u}}{\omega_{0}}$ is the term appearing in $\mathcal{F}_{\omega_{0}}$. This property can be deduced from the residue computation in the local coordinates in (58) which turns into the leading order term of $\mathcal{F}_{\omega_{0}}(\tilde{u},g)$ at $g=0$. In fact, at any order contributing to (71), we encounter such integrals which generate the corrections to $\log\mathcal{C}^{\mathrm{i}}_{0}$ and $\log\mathcal{C}^{\mathrm{ii}}_{0}$ terms in (73) and (74) respectively. Eventually, their total contribution to $e^{-W_{0,1}}$ becomes
$e^{-W_{0,1}}=e^{-\mathcal{G}(\tilde{u},g)}e^{-2\mathcal{F}(\tilde{u},g)\log\mathcal{C}^{\mathrm{i}}}\,,$ (75)
@@
$e^{-W_{0,1}}=e^{-\mathcal{G}(\tilde{u},g)}e^{-\mathcal{F}(\tilde{u},g)\log\mathcal{C}^{\mathrm{ii}}}\,,$ (76)
@@
In this form, the computations of $W_{0,1}$ can be carried out by using Mellin transformation at each order as described in DP1 ; Zinn-Justin:2004qzw ; Kamata:2021jrs . Then, following Zinn-Justin:2004qzw , we introduce normalizations at the double turning points. Then, dropping $\omega_{0}$ subscripts, we write the final generic expressions for $\Pi_{\mathrm{NP}}^{\mathrm{i}}$ and $\Pi_{\mathrm{NP}}^{\mathrm{ii}}$ as
$\tilde{\Pi}^{\mathrm{i}}_{\mathrm{NP}}$ $=e^{-\mathcal{G}(\tilde{u},g)}\,\frac{2\pi\left(\frac{g}{\mathcal{C}^{\mathrm{i}}}\right)^{-2\mathcal{F}(\tilde{u},g)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}(\tilde{u},g)\right)\right]^{2}}\,,\qquad\tilde{\Pi}^{\mathrm{ii}}_{\mathrm{NP}}=e^{-\mathcal{G}(\tilde{u},g)}\,\frac{\sqrt{2\pi}\,\left(\frac{g}{\mathcal{C}^{\mathrm{ii}}}\right)^{-\mathcal{F}(\tilde{u},g)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}(\tilde{u},g)\right)}\,,$ (79)
@@
$\mathcal{C}^{\mathrm{i}}$ $=2\,\omega_{0}\left(x_{1}-x_{0}\right)^{2}\,\mathrm{exp}\left\{\int_{x_{0}}^{x_{1}}{\mathrm{d}}x\,\left[\frac{\omega_{0}}{\sqrt{2\tilde{V}(x)}}-\frac{1}{x-x_{0}}-\frac{1}{x_{1}-x}\right]\right\}\,,$ (80)
$\mathcal{C}^{\mathrm{ii}}$ $=2\,\omega_{0}\,\left(x_{1}-x_{0}\right)^{2}\mathrm{exp}\left\{2\int_{x_{0}}^{x_{1}}{\mathrm{d}}x\,\left[\frac{\omega_{0}}{\sqrt{2\tilde{V}(x)}}-\frac{1}{x_{1}-x}\right]\right\}\,.$ (81)
Note that we inserted the prefactor $\omega_{0}$, which originally appears in $\frac{g}{\omega_{0}}$ terms in (70) in the definitions of $\mathcal{C}^{\mathrm{i}}$ and $\mathcal{C}^{\mathrm{ii}}$ for convenience.
These expressions recover ZJJ’s results in Zinn-Justin:2004vcw ; Zinn-Justin:2004qzw for the bottom of wells and generalizes them for all types of locally harmonic saddles. At the barrier tops, $\omega_{0}$ is an imaginary constant and another imaginary contribution arises from $\sqrt{2\tilde{V}(x)}$. As a result, the term $\frac{\omega_{0}}{\sqrt{2\tilde{V}(x)}}$ is real and only possible imaginary contribution is due to $\omega_{0}$ term. Moreover, the log divergence(s) at double turning point(s) is(are) cancelled by the normalization terms in (80) and (81).